OpenStudy (anonymous):

What is the x-intercept, y-intercept, Rewrite the equation in its slope-intercept form, and what is the slope for: i(-2,-3), ii (6,3), iii(1,-3/4) , iv(-6,-6)

OpenStudy (anonymous):

x-intercept is a point on the graph when y is zero y-intercepts is a point on the graph when x is zero for example: find the x-intercepts and y-intercepts from this equation :

OpenStudy (anonymous):

for example : \[25x ^{2}+4y ^{2}=9\] x-intercepts : y=0 \[25x ^{2}+4(0)^{2}=9\] \[x ^{2}=9/25\] \[x=\pm 3/5\] so x-intercepts at the point (3/5,0) and (-3.5,0)

OpenStudy (anonymous):

y-intercepts : x=0 \[25(0)^{2}+4y ^{2}=9\] \[y ^{2}=9/4\] \[y=\pm 3/2\] so Y-intercepts at point (0,3/2) and (0,-3/2)

OpenStudy (anonymous):

slope for i (-2,-3) and ii (6,3) \[(y-y1)/(y2-y1)=(x-x1)/(x2-x1)\] \[y=1/2x-6\] standard formula : \[y=mx+b\] so the slope is m=1/2